HUBBLE SPACE TELESCOPE DAILY REPORT #2300 PERIOD COVERED: 0000Z (UTC) 02/02/99 - 0000Z (UTC) 02/03/99 Daily Status Report as of 034/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.3 Completed STIS/CCD/WF/PC-2 7367 (Multi-Wavelength Imaging of Circumstellar Disks in the Orion Nebula) The Space Telescope Imaging Spectrograph (CCD) and the WF/PC-2 were used to make multi-wavelength imaging of circumstellar disks (NGC-1976) seen as silhouettes against the bright emission-line background in the Orion Nebula. With these data one should be able to examine the radial structure of the disks in great detail, and over a wide spatial extent. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.5 Completed WF/PC-2 7447 (The Cosmological Evolution of Quasar Host Galaxies) The WF/PC-2 was used to make observations of galaxy B2-0110+29 as part of a survey to study the cosmological evolution of galaxies which host quasars. Other point spread function observations were taken to test the calibration of the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD/MA1/MA2 7389 (Chromospheres and Transition Layers in Hyades and Pleiades F stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the F class star HD27808 which is located in the Hyades star cluster. This proposal seeks to obtain additional empirical constraints on the heating mechanisms for stellar chromospheres, transition layers and coronae. A wavelength calibration obseration was also taken. About two hours into this two-hour, 45-minute proposal, while the high voltage was on, the MCE-1 reset described in yesterday's report and HSTAR 6964 occurred. Observations were affected. Otherwise, the proposal completed as planned, and no further problems were reported. 1.7 Completed Two Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.9 Completed WF/PC-2 6360 (A Snapshot Survey of "Radio Loud" IRAS Galaxies) The WF/PC-2 was used to make snapshot survey observations of the radio loud IRAS galaxy GAL-1747+4. The WF/PC-2 was used to discern high surface brightness structures in its post-starburst active nuclear regions. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed STIS/CCD 7403 (Measuring Black Hole Masses in Broad-Lined AGNs) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of broad-lined active galactic nuclei (AGN) located in galaxy NGC-3245. Broad emission lines in AGNs are considered one of the most direct signatures of a large central mass, which is presumed to be a black hole. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.11 Completed STIS/CCD 7354 (Kinematics of Gaseous Disks in the Nuclei of Liners with Compact Flat Spectrum Radio Sources) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclei of the liner in the compact flat spectrum source galaxy NGC-5635. Many active galactic nuclei channel energetic particles into jets that feed powerful extended lobes of radio emission; others remain quite compact. Both classes of active galaxies are suspected to be powered gravitationally by the accretion of surrounding material onto a massive black hole. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.12 Completed Two Sets of WF/PC-2 6358 (UV Imaging of Circumnuclear Starburst Rings) The WF/PC-2 was used to make ultraviolet snapshot observations of galaxies NGC-4995 and NGC-4321 in order to study their circumnuclear starburst rings. A homogeneous database of properties of the star-forming region will be generated. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared twice. A TTR was written for a data hit and 142 missing minor frames that occurred at 033/2137Z. As documented in HSTAR 6967, the STIS MCE-2 reset at 034/06/1828Z while the low voltage was on and while in an SAA contour. MAMA-2 recovery will be accomplished at 034/143900Z, the time of the next high voltage on commanding via the SMS. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH